OpenStudy (anonymous):

Part one:Solve the inequality -5x + 8 less than or greater to x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (texaschic101):

-5x + 8 <= x - 4 --- subtract x from both sides -5x - x + 8 <= x - x - 4 --- combine like terms -6x + 8 <= -4 --- subtract 8 from both sides -6x + 8 - 8 <= -4 - 8 --- combine like terms -6x <= - 12 --- divide both sides by -6, but remember that when you divide/multiply by a negative, the sign changes. x >= 2 Leatn this : if there is an equal sign, the circle is closed no equal sign, is an open circle < (less then) means shading to the left >(greater then) means shading to the right so your graph would be : closed circle on -2, shading to the right

OpenStudy (anonymous):

thank you!!